# Galley: Modern Query Optimization for Sparse Tensor Programs

### v3

@@ 3.1. Tensor Index Notation
The input to our system, Galley, is an extended version of Einsum notation which we call tensor index notation. Traditional Einsum notation allows for a single summation wrapped around a multiplication operation. For instance, you can describe triangle counting in a graph where $E_{ij}$ is the adjacency matrix with the following Einsum statement,
@@
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting of aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether they occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$ which is $1$ if $i$ occurs in at least one triangle and 00 otherwise.
@@
Tensor compilers like Halide, TACO, and Finch each build off of similar core notations, adding additional structures like FOR-loops to allow users to specify algorithmic choices(Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Ahrens et al., 2024). Crucially, the vast majority of operations in array programming frameworks like Numpy can be expressed as operations in tensor index notation So, while we focus on this notation in this paper, traditional tensor workflows can be captured and optimized in this framework.
Figure 3. Galley Overview
@@ 3.2. Sparse Tensor Compilers
In the last decade, the compilers community has developed a series of sparse tensor compilers and shown that they can produce highly efficient code for sparse tensor computations. We use this work as our execution engine, so we take a moment here to explain their important concepts.
@@
There are many different ways to represent sparse tensors, and the best approach depends on the data distribution and the workload. Work in this space has converged on the fibertree abstraction for describing the space of formats(Kjolstad et al., 2017; Sze et al., 2017). In this formalism, a tensor format is a nested data structure like the one in Fig. 2. Each layer stores the non-fill (e.g. non-zero) indices in a particular dimension, conditioned on the earlier dimensions, and pointers to the next dimension’s non-fill indices. These layers can be represented in any format that allows for iteration and lookup. In this work, we consider sorted lists, hash tables, bytemaps, and dense vectors, and they each have different performance characteristics in terms of iteration, lookup, and memory footprint. For example, compressed sparse row (CSR) is a common format for sparse matrices. It stores the row dimension as dense vector where each entry points to the set of non-zero columns for that particular row. This set of non-zero columns is then stored in a sorted list, i.e. in a compressed sparse format. Importantly, this abstraction requires tensors to be accessed in the order that they are stored (e.g. row-then-column in the case of CSR) which restricts the set of valid loop orders as we describe next.
@@ C.   Physical Plan
Figure 4. Query Plan Dialects
@@
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL) which consists of for-loops, in-place aggregates (e.g. +=+=+ =), and arithmetic over indexed tensors (e.g. $A[i,j]*B[j,k]$). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 00 is the annihilator of multiplication (i.e. $x*0=0$), only the values of $i$ which map to non-zero entries in $X$ and $\theta$ need to be processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$. Considering the illustration in Fig. 2, we can see how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices which are non-zero in $X[i,\_]$, i.e. the non-zero columns that occur in each row.
@@
Once the compiler has determined which tensors’ non-zero indices need to be merged to iterate over a particular index, there are several algorithms it can apply. All formats allow for both ordered iteration and lookup operations, so one algorithm is to iterate through the indices of all inputs, similar to a merge join. This is highly efficient per operation. However, it’s linear in the total size of all inputs, even if one is much smaller than the others. Another method is to iterate through one input’s level and lookup that index in the others. In this work, we generally take the latter approach, as described in Sec. 6.3. We refer to the mode of an individual tensor (such as “iterate” or “lookup”) as an access protocol and the overall strategy as a merge algorithm(Ahrens et al., 2023).
@@ 4. Galley Overview
In this section, we provide a birds-eye view of Galley and show how an input program is transformed, first to a logical plan then to a physical plan then to an STC program, as illustrated in Fig.  3. The first transformation from input program to logical plan is handled by Galley’s logical optimizer, which extends the variable elimination framework to break the program down into a sequence of aggregations. The second transformation, from logical plan to physical plan, is performed by Galley’s physical optimizer, which determines the loop order, tensor formats, and merge algorithms for each logical query. These steps are each represented by a dialect of our query plan language whose grammar is defined in Fig. 4. In the following discussion, we use this grammar as a guide and show how our example program, logistic regression, would be transformed through these steps.
The input program dialect is equivalent to the tensor index notation defined in Sec. 3.1. Pointwise functions like $A_{ij}*B_{jk}$ are represented with Map. Aggregates like $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Our logistic regression example from Eq. (1) is defined in this dialect as follows,
@@
Note that this notation is compatible with array APIs like Numpy which don’t have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices. Also, note that aggregates can be over multiple indices, e.g. Agg(+, i, j, k, ...).
@@ 4.1. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, is to break down the input program into a sequence of simple aggregates. This is enforced by converting the input program to a logical plan. This dialect is a restriction of the input dialect where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query which produces a semantically equivalent output. To do this efficiently, it must minimize the total cost of all queries in the logical plan.
As an example, our logistic regression program above can be translated into the following logical plan,
@@
In this plan, the first query isolates the sum over the $j$ index while the second query performs the remaining sigmoid operation on the result. Note that the latter query uses a no-op aggregate to represent an element-wise operation while conforming to the logical dialect.
@@ 4.2. Physical Plan
Given the logical plan, Galley’s physical optimizer determines the implementation details needed to convert each logical query to a STC kernel. Specifically, it defines the loop order of each compiled kernel, the format of each output, and the merge algorithm for each index. As above, this is expressed by converting the logical plan to a physical plan described in the most verbose and constrained dialect. To avoid out-of-order accesses, we require that the index order of inputs and aliases are concordant with the loop order, so the physical optimizer may insert additional queries to transpose inputs. Therefore, each logical query corresponds to one or more physical queries.
Using this language, we can precisely express the program from Fig. 1 as follows where it means iterate and lu means lookup.
@@
The first query computes the sum by iterating over the valid i indices for X, iterating over the j indices in the intersection of X[i,_] and $\theta$, and accumulating their product in a dense vector over the i indices. The second query runs over this output and applies the sigmoid function, returning the result as a dense vector.
@@ 4.3. Execution
Once Galley has generated a physical plan, the execution is very simple. For each physical query, it first translates the expression into an STC kernel definition and calls the STC to compile it. Then, Galley injects the tensors referenced by inputs and aliases and executes the kernel, storing the resulting tensor in a dictionary by name. After all queries have been computed, we return the tensors requested in the input program by looking them up in this dictionary. In Sec. 8, we describe additional optimizations that improve this workflow by doing just-in-time physical optimization and common sub-expression elimination.
@@ 5. Logical Optimizer
Given the plan dialects above, we now describe the logical optimizer, which takes in an input program and outputs a semantically equivalent logical plan. Specifically, the logical optimizer converts each query in the input program to a sequence of logical queries where the last one produces the same output as the input query. There are many valid plans, and the optimizer searches this space to identify the cheapest one. In this section, we briefly define ”cheapest” in this context before outlining the complex space of valid logical plans that are considered. Lastly, we explain the algorithms that we use to perform this search.
5.1. Canonicalization & Pointwise Distributivity
The first step in our logical optimization is to canonicalize the input program with a few simple rules that we apply exhaustively; 1) merging nested Map operators 2) merging nested Agg operators 3) lifting Agg operators above Map operators when possible 4) renaming indices to ensure uniqueness. This compresses the input program and makes our reasoning simpler later by ensuring that operator boundaries are semantically meaningful.
Next, we consider whether to distribute point-wise expressions. This may or may not result in a better plan because it both makes operations more sparse and produces a larger expression.
@@ 5.1. Canonicalization & Pointwise Distributivity
Consider the following expression and its distributed form,
@@
If all inputs are dense, the non-distributed form is more efficient because it results in fewer terms and has the same computational cost per term. However, if $X_{ij}$ is sparse and $U_{i},V_{j}$ are dense, then the distributed form could be more efficient because all terms can be computed in time linear w.r.t. the sparsity of $X_{ij}$. Note that the squaring operation here is a point-wise function not a matrix multiplication.
Because apply pointwise distributivity can produce asymptotic performance improvements, it cannot be ignored. To address this, we perform logical optimization for both the distributed expression and the original and use the cheaper one.
@@ 5.2. Cost Model
Overall, Galley’s logical optimizer attempts to minimize the time required to execute the logical program. Because logical queries do not correspond to concrete implementations, our logical cost model aims to approximate this time without reference to the particular implementation that the physical optimizer will eventually decide on. This approximation considers on two factors: (1) the number of non-fill entries in the output tensor and (2) the amount of computation (i.e. the number of FLOPs) needed to produce the output. The former corresponds to the size of the tensor represented by Agg, $nnz(\texttt{Agg})$, and the latter corresponds to the size of the tensor represented by the MapExpr within, $nnz(\texttt{MapExpr})$. We assume that the inputs are in memory, hence there is no cost for reading inputs from disk. We then performed a simple regression to associate each of these costs with a constant, and we add them to produce our overall cost, $c$,
@@
To estimate $nnz(\texttt{Agg})$ and $nnz(\texttt{MapExpr})$, we introduce a sparsity estimation framework in Sec. 7.
@@ 5.3. Variable Elimination
The core of our logical optimizer is the variable elimination (VE) framework. With this view, a logical plan for an input query is defined by an order on the indices being aggregated over, i.e. an elimination order. Given this order, we can construct a valid logical plan in the following way. Iterating through the elimination order one index at a time, we (1) identify the minimal sub-query needed to aggregate it out of the expression, (2) create a new logical query representing the result of that sub-query, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
@@
Consider optimizing the following matrix chain multiplication,
@@
The elimination order $jkl$ corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries,
@@
Unlike traditional VE over sum-product queries, we need to handle complex trees of pointwise operators and aggregates. This makes the identification of minimal sub-queries challenging as we need to carefully examine the expression’s algebraic properties. Given a strategy for this, the core problem of optimizing VE is to search the space of elimination orders for the most efficient one. In the worst case, this takes exponential time w.r.t. the number of indices being aggregated over. In the following sections, we describe 1) how we identify minimal sub-queries and 2) our search algorithm for finding the optimal elimination order.
5.4. Identifying Minimal Sub-Queries
In this section, we show how to identify the minimal sub-queries (MSQ) needed to aggregate over an index. In sum-product expressions, the MSQ is just the sum over that index wrapping the product of the tensors that are indexed by it. However, it’s not immediately clear how to do this for more complex input programs which include an arbitrary mix of operators and aggregates. Fortunately, we show that identifying the MSQ corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of the operation at each node to determine how to proceed.
Annotated Expression Tree The annotated expression tree (AET) is constructed by examining the nested structure of Agg, Map, Input, and Alias nodes in the input query. We start by removing all Agg nodes and annotating their inner expressions with (Idx, Op) for each index being aggregated over. We then replace all Map nodes with their operator to get the final tree where every internal node is a function and every leaf is either an Input or an Alias.
Figure 5. Annotated expression tree for logistic regression, $\sigma(\sum_{jpc}(S_{ipc}P_{pj}\theta_{j}+S_{ipc}C_{cj}\theta_{j})$

### v4

@@ 3.1. Tensor Index Notation
Input to Galley is written in an extended version of Einstein Summation (Einsum) notation that we call tensor index notation(Albert et al., 1916). Traditional Einsum notation permits a single summation wrapped around a multiplication. For instance, you can describe triangle counting in a graph with adjacency matrix $E_{ij}$ using the following statement:
@@
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether the entities occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$, which is $1$ if $i$ occurs in at least one triangle and 00. This can be written in tensor index notation as:
@@
Tensor compilers like Halide, TACO, and Finch each build off of similar core notations, adding additional structures like FOR-loops to let users specify algorithmic choices (Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Ahrens et al., 2024). Crucially, the vast majority of operations in array programming frameworks like Numpy can be expressed as operations in tensor index notation. Therefore, though we focus here on this notation, traditional tensor workflows can be captured and optimized in this framework.
Figure 2. Galley overview.
@@ 3.2. Sparse Tensor Compilers
Over the last decade, compiler researchers have developed a series of sparse tensor compilers and shown that they produce highly efficient code for sparse tensor computations(Kjolstad et al., 2017; Ahrens et al., 2024). We use this work as our execution engine, so we briefly explain its important concepts below.
Figure 3. Fibertree format abstraction.
@@
There are many different ways to represent sparse tensors, and the optimal approach depends on the data distribution and the workload. Work in this space has converged on the fibertree abstraction for describing the space of formats (Kjolstad et al., 2017; Sze et al., 2017). In this formalism, a tensor format is a nested data structure resembling the one in Fig. 3. Each layer stores the non-fill (e.g., non-zero) indices in a particular dimension, conditioned on earlier dimensions, and pointers to the next dimension’s non-fill indices. These layers can be represented in any format that enables iteration and lookup.
In this work, we consider sorted lists, hash tables, bytemaps, and dense vectors, which each perform differently in terms of iteration, lookup, and memory footprint. For example, the compressed sparse row (CSR) is a common format for sparse matrices. It stores the row dimension as a dense vector, where each entry points to the set of non-zero columns for that particular row. This set of non-zero columns is then stored in a sorted list, i.e., in a compressed sparse format. Importantly, this abstraction requires tensors to be accessed in the order in which they are stored (e.g., row-then-column in the case of CSR), which restricts the set of valid loop orders, as we describe next.
@@ C.   Physical Plan
Figure 4. Query plan dialects.
@@
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL); it consists of for-loops, in-place aggregates (e.g., +=+=+ =), and arithmetic over indexed tensors (e.g., $A[i,j]*B[j,k]$). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 00 is the annihilator of multiplication (i.e., $x*0=0$), only the values of $i$ that map to non-zero entries in $X$ and $\theta$ are processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$; Fig. 3 shows how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices that are non-zero in $X[i,\_]$, i.e., the non-zero columns that occur in each row.
@@
Once the compiler has determined which tensors’ non-zero indices must be merged to iterate over a particular index, it can apply several algorithms. All formats enable both ordered iteration and lookup operations; therefore, one algorithm iterates through the indices of all inputs, similar to a merge join, which is highly efficient per operation. However, this algorithm is linear in the total size of all inputs even if one is much smaller than the others. Another method is to iterate through a single input’s level and lookup that index in the others. In this work, we take the latter approach, as described in Sec. 6.3. We refer to the mode of an individual tensor (such as “iterate” or “lookup”) as an access protocol and the overall strategy as a merge algorithm (Ahrens et al., 2023).
@@ 4. Galley Overview
We now provide a high-level view of Galley. We show how it transforms an input program to a logical plan then to a physical plan that is executed by an STC, as illustrated in Fig.  2. These steps are each represented by a dialect of our query plan language, whose grammar is defined in Fig. 4. In the following discussion, we use this grammar as a guide to show how our example program, i.e., logistic regression, would be transformed through these steps.
The input program dialect is equivalent to the tensor index notation defined in Sec. 3.1. Pointwise functions such as $A_{ij}*B_{jk}$ are represented with Map. Aggregates such as $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Our logistic regression example from Eq. (1) is defined in this dialect as
@@
Note that this notation is compatible with array APIs like Numpy that do not have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices. Also, note that aggregates can be over multiple indices, e.g., Agg(+, i, j, k, ...).
@@ 4.1. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, breaks down the input program into a sequence of simple aggregates. This is enforced by converting the input program ( 4.A) to a logical plan ( 4.B). This dialect is a restriction of the input dialect, where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query, which produces a semantically equivalent output. To do this efficiently, it must minimize the total cost of all queries in the logical plan.
Our logistic regression program above is not a valid logical plan because the outer expression is a pointwise function not an aggregate. However, it can be translated into the following logical plan
@@
In this plan, the first query isolates the sum over the $j$ index, while the second query performs the remaining sigmoid operation on the result. Note that the latter query uses a no-op aggregate to represent an element-wise operation while conforming to the logical dialect.
@@ 4.2. Physical Plan
Given the logical plan, Galley’s physical optimizer determines the implementation details needed to convert each logical query to an STC kernel. Specifically, it defines the loop order of each compiled kernel, the format of each output, and the merge algorithm for each index. As above, this is expressed by converting the logical plan to a physical plan described in the most constrained dialect. To avoid out-of-order accesses, we require that the index order of inputs and aliases are concordant with the loop order, so the physical optimizer may insert additional queries to transpose inputs. Therefore, each logical query corresponds to one or more physical queries.
Using this language, we can precisely express the program from Fig. 1 as follows, where it means iterate and lu means lookup.
@@
The first query computes the sum by iterating over the valid i indices for X, iterating over the j indices in the intersection of X[i,_] and $\theta$, and materializing (hence Mat) their product in a dense vector over the i indices. The second query runs over this output and applies the sigmoid function, returning the result as a dense vector.
@@ 4.3. Execution
Once Galley has generated a physical plan, the execution is very simple. For each physical query, it first translates the expression into an STC kernel definition and calls the STC to compile it. Then, Galley injects the tensors referenced by inputs and aliases and executes the kernel, storing the resulting tensor in a dictionary by name. After all queries have been computed, it returns the tensors requested in the input program by looking them up in this dictionary.
@@ 5. Logical Optimizer
Given the plan dialects above, we now describe the logical optimizer, which receives an input program (Dialect  4.A) and outputs a semantically equivalent logical plan (Dialect  4.B). Specifically, the logical optimizer converts each query in the input program to a sequence of logical queries, where the last query produces the same output as the input query. There are many valid plans, and the optimizer searches this space to identify the cheapest one. We now briefly define ”cheapest” in this context before outlining the complex space of logical plans that are considered. Finally, we explain the algorithms that we use to perform this search.
5.1. Normalization & Pointwise Distributivity
The first step in logical optimization is to normalize the input program with a few simple rules that we apply exhaustively: (1) merge nested Map operators, (2) merge nested Agg operators, (3) lift Agg operators above Map operators, when possible, and (4) rename indices to ensure uniqueness. Applying these rules compresses the input program and makes our reasoning simpler in later steps by ensuring that operator boundaries are semantically meaningful.
Next, we consider whether to distribute pointwise expressions. Doing so may or may not yield a better plan because it both makes operations more sparse and produces larger expressions.
@@ 5.1. Normalization & Pointwise Distributivity
Consider the following expression which computes the loss function for the alternating least squares (ALS) algorithm and its distributed form:
@@
If all inputs are dense, the non-distributed form is more efficient because it results in fewer terms and has the same computational cost per term. However, if $X_{ij}$ is sparse and $U_{i},V_{j}$ are dense, then the distributed form is more efficient because all terms can be computed in time linear w.r.t. the sparsity of $X_{ij}$. Note that the squaring operation here is a pointwise function, not a matrix multiplication.
To take advantage of this potentially asymptotic performance improvement, Galley performs a greedy search for the optimally distributed expression. At each step, it considers all single applications of distributivity in the expression. It then runs variable elimination for each (described later in this section) and computes the cost an optimal logical plan. If applying distributivity improved on the cost of the original expression, it continues. If not, it returns the optimal logical plan discovered so far. Lastly, we additionally consider the expression derived from applying distributivity exhaustively.
@@ 5.2. Cost Model
Overall, Galley’s logical optimizer attempts to minimize the time required to execute the logical program. Because logical queries do not correspond to concrete implementations, our logical cost model aims to approximate this time without reference to the particular implementation that the physical optimizer will eventually decide on. This approximation considers two factors: (1) the number of non-fill entries in the output tensor and (2) the amount of computation (i.e., the number of FLOPs) needed to produce the output. The former corresponds to the size of the tensor represented by Agg, $nnz(\texttt{Agg})$, and the latter corresponds to the tensor size represented by the MapExpr within, $nnz(\texttt{MapExpr})$. We assume that the inputs are in memory; hence, there is no cost for reading inputs from disk. We then perform a simple regression to associate each cost with a constant, and we add them to produce our overall cost, $c$, as follows:
@@
To estimate $nnz(\texttt{Agg})$ and $nnz(\texttt{MapExpr})$, we use the sparsity estimation framework described in Sec. 7.
@@ 5.3. Variable Elimination
The core of our logical optimizer is the variable elimination (VE) framework. With this view, the logical plan for an input query is defined by an order on the indices being aggregated over, i.e., an elimination order. Given this order, we can construct a valid logical plan by iterating through the elimination order one index at a time in order to (1) identify the minimal sub-expression needed to aggregate over it, (2) create a new logical query representing the result of that sub-expression, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
@@
Consider optimizing the following matrix chain multiplication:
@@
The elimination order $jkl$ corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries:
@@
Unlike traditional VE for sum-product queries, we support complex trees of pointwise operators and aggregates. This makes identifying minimal sub-queries challenging since we must carefully examine the expression’s algebraic properties. Given a strategy for this, the core problem of optimizing VE is to search the space of elimination orders for the most efficient one. In the worst case, this takes exponential time w.r.t. the number of indices being aggregated over. In the following sections, we describe how we identify minimal sub-queries and our search algorithm for finding the optimal elimination order.
5.4. Identifying Minimal Sub-Expressions
We now explain how to identify the minimal sub-expressions (MSEs) needed to eliminate an index. In sum-product expressions, the MSE is simply a summation over that index wrapping the product of the tensors that are indexed by it. However, it is not immediately clear how to do this for more complex input programs. Fortunately, we show that identifying MSEs corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of the operation at each node to determine how to proceed.
Figure 5. Annotated expression tree for logistic regression
$\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
Annotated Expression Tree. The annotated expression tree (AET) is constructed by examining the nested structure of Agg, Map, Input, and Alias nodes in the input query. To do this, Galley first removes all Agg nodes and annotates their inner expressions with (Idx, Op). It then replaces all Map nodes with their operator to get the final tree, where every internal node is a pointwise function and every leaf is either an Input or an Alias.